# V2X Cooperative Perception for Autonomous Driving: Recent Advances and Challenges

### v4

@@ III Development of Cooperative Perception: From Early Time Exploration till Recent Research
CP involves the exchange of detected obstacles or perception data among agents, thereby expanding their perceptual range, enhancing situational awareness, and ensuring safety [101]. In the context of driving automation, Cooperative Prediction entails receiving intention or desired trajectory information from neighbouring vehicles, thereby enhancing the accuracy and efficiency of motion prediction. This approach offers the potential to improve strategic ego-planning functionality [102].
In its nascent phase, research in cooperative transportation systems primarily centred on devising statistical signal processing and rule-based algorithms, often supplemented by data from artificial sources or other sensors such as radar and LiDAR. For example, a study outlined in [103] introduced a particle filter-based multi-target tracking framework that expanded detection range by utilizing artificial radar measurements and facilitated vehicle position sharing via V2V communication. Simulation outcomes underscored CP’s potential to enhance the performance of intelligent vehicle systems by amalgamating information from multiple vehicles, thus complementing single-vehicle detection systems. Similarly, in [104], research aimed to construct a synthesized representation of the surrounding environment for multi-vehicle CP. This was achieved by transforming the visual perspectives of two vehicles into a shared spatial reference system predicated on the estimated relative positions between the two vehicles. Subsequently, a laser line was employed to generate an approximate depth map of visual perceptions, establishing correspondence between the images. This innovative algorithm conferred the ego vehicle with the ability to perceive objects obstructed in the foreground. Nevertheless, it had notable limitations, primarily stemming from imprecise pose estimation and the use of an approximate depth map.
CP can be viewed as a mechanism designed to tackle the challenge of map merging, aligning perception data across multiple agents, and integrating it into a consistent reference frame [105]. Achieving precise results in this context necessitates temporal and spatial alignment before the post-data fusion phase. Early studies addressed this challenge by employing motion models and estimation algorithms rooted in statistical signal processing. In [105], researchers leveraged a Car2X-based perception module for object-level data fusion, utilizing an Unscented Kalman Filter to align the received bounding box proposals from communicating vehicles and roadside stations. This module facilitated the prediction of the object’s state directly to the current time step, utilizing the Constant Turn Rate and Acceleration motion model within the spatial coordinate frame of the objects. Similarly, in the realm of cooperative tracking for multiple vehicles, the study presented in [106] extended a Gaussian Mixture Probability Hypothesis Density (GM-PHD) filter with a collaborative fusion algorithm. State propagation relied on the Constant Turn Rate and Velocity motion model, accommodating variable delays in wireless communication by condensing the communicating tracks into GM-PHD densities. It is noteworthy, however, that these motion models, while effective in their own right, impose significant constraints that limit their applicability within the domain of vehicular motion.
The rule-based algorithm, devoid of a learning process, offers a computationally efficient approach with demonstrated engineering viability. In [107], an engineering-focused multi-modal CP system was introduced to facilitate cooperative autonomous driving. This system employed a mirror-neuron-inspired intention awareness algorithm. The mirror neuron concept refers to neurons that activate when an organism performs or observes a similar action in others, forming the basis for generating intention awareness. Specifically, it enabled all vehicles’ collective synthesis of anticipatory trajectories by simulating drivers’ behaviours. However, it is essential to note that this approach relied heavily on precise estimation outcomes from each agent within the vehicular network due to its use of track-level data fusion, potentially affecting model effectiveness.
@@
Many conventional methods face challenges in effectively extracting attributes from shared information that accurately represent the necessary entities for fulfilling the requirements of driver assistance systems. In recent years, deep learning technology has demonstrated significant advancements in single-vehicle perception. As a result, numerous researchers have been actively involved in developing learning-based CP frameworks to overcome the limitations of traditional approaches. For instance, in the realm of cooperative object detection and classification, researchers have leveraged advanced deep learning techniques such as the YOLO network [108] and DenseNet [109] to extract 3D information from shared objects [110]. Further discussion on recent research in this area will be presented in Sections IV and V.
Traditional approaches in CP have typically relied on late collaboration strategies due to the complexity of feature extraction and limited communication capabilities. With the rapid progress of deep learning, intermediate collaboration has gained prominence in multi-agent cooperation, primarily due to its ability to alleviate the burden on communication resources [111]. Nevertheless, it risks introducing extraneous and unnecessary information, particularly in dynamic networking conditions. To address this challenge, optimizing cooperative information and collaborator selection becomes imperative within a CP system. Furthermore, deep representation introduces higher levels of abstraction and implicitness in the spatial interactions between agents, necessitating suitable data alignment and information fusion algorithms. Developing a unified CP framework to address these issues is of significant importance in advancing Cooperative Driving Automation (CDA).
@@ IV-A Framework Overview
TABLE III: Summary of the Methods of Preparation for Transmission.
@@
CAS Who2com [115] 2020 -- Camera AirSim-CP Intermediate Segmentation $\surd$
When2com [116] 2020 -- Camera AirSim-MAP Intermediate Segmentation $\surd$
@@
PIS DyDisse [123] 2019 V2X Camera/LiDAR -- Late Tracking $\times$
VehIdenti [124] 2023 V2X -- CARTERY -- -- $\times$
RedunMitig [125] 2019 V2X -- SUMO Late Detection $\times$
Look-ahead [126] 2019 V2X -- SUMO Late Detection $\times$
AICP [127] 2022 V2V Camera -- Late Detection,Tracking $\surd$
@@
DRLCP [130] 2020 V2X Camera CIVS Late -- $\times$
@@ IV-B Preparation for Transmission
In CP systems, continuous information exchange among multiple agents is crucial for achieving comprehensive traffic awareness. However, real-world application scenarios often face limitations in communication bandwidth. Excessive network load increases the risk of important data packets being delayed or lost, which could compromise the performance and safety of CP systems. Therefore, it is essential to balance the trade-off between perception performance and communication efficiency. To address this challenge, researchers have proposed various strategies aimed at reducing the volume of shared data prior to transmission and alleviating the burden on vehicular networks. As summarized in Table III, these strategies can be grouped into four main approaches:
@@ IV-B1 Cooperative Agent Selection
The attention mechanism, recognized for its efficacy in improving model efficiency by focusing on relevant data, offers promise in selecting suitable agents for the ego CAV. In this context, a multi-stage handshake communication mechanism was proposed in Who2com [115], employing the general attention mechanism [152] to select the appropriate collaborators. This mechanism computes matching scores among agents, selectively engaging the most pertinent agents to optimize bandwidth usage. Based on Who2com, When2com [116] introduced scaled general attention to ascertain optimal communication timings, enabling the ego agent to initiate communication solely under conditions of insufficient information, thereby conserving network resources efficiently. Although this multi-stage communication process proves effective, its implementation and optimization in dynamically varying environments pose significant challenges. To address these challenges, [117] proposed a dynamic communication mechanism that allows vehicles to dynamically determine which infrastructures to communicate with based on the learned attention weights, thus striking a balance between detection accuracy and communication spectrum resources.
@@
Cooperative agent selection can also be likened to a Multi-agent Path Finding (MAPF) problem, which plans paths for multiple agents [155]. Inspired by MAPF, a selective communication algorithm [120] was proposed to determine the subset of cooperative vehicles based on information gains from connected vehicles. The ego-vehicle estimates information gains by comparing its detected objects’ positions to received 2D centre positions from other connected vehicles. The vehicles associated with the highest information gain are selected to form the communication scope. In contrast to random selection, this algorithm achieved higher success rates with minimal additional communication costs.
All the aforementioned methods primarily focused on spatial correlations, overlooking the significance of temporal dependencies in collaborator selection. To address this gap, IoSI-CP [121] capitalized on the importance of semantic information from both temporal and spatial dimensions to select optimal collaborators. This approach leveraged a Graph Neural Network (GNN) to capture inherent correlations between the ego vehicle and its neighbouring agents. Agents with weights of zero or lower were considered ineligible for collaboration, which eliminates unsuitable collaborators responsible for noise. Nevertheless, this approach relied on one-dimensional weights that may not fully capture the intricacies addressed by high-dimensional representation methods. InterCoop [122] computes an interaction score for each vehicle by analyzing the geometric properties of the road network in conjunction with the temporal dynamics of the trajectories, prioritizing those most relevant to the prevailing driving conditions.
@@ IV-B2 Perception Information Selection
The rule-based approach to CPM generation/selection entails designing rules that dictate which aspects of detected object information should be included in a CPM. For example, the work in [123] proposed a rule that transmits information about an object’s mode only if the object is turning or accelerating. Similarly, in [124], visual information from detected vehicles was used to determine their communication IDs, which are then used to integrate CPMs with local perception information. All detected vehicles that have exchanged CAM messages will not be considered in future CPMs. Standardization efforts, led by the European Telecommunications Standards Institute (ETSI), have also introduced CPM generation rules, specifying what information CPMs should include and when to generate new messages. The work in [125] and [126] has proposed improved algorithms for these rules to filter out redundant information about detected objects that nearby vehicles have recently transmitted. While rule-based methods are straightforward to implement, they rely primarily on changes in the position and speed of detected objects.
@@
Distance-based methods prioritize data based on distance metrics, considering the spatial location relative to the ego vehicle’s onboard sensors to measure data point significance. For instance, the Augmented Informative Collaborative Perception (AICP) system [127] employed a fitness sorting algorithm based on Mahalanobis distance to select crucial information at the application level, effectively filtering out less relevant packets with minimal delay. In a separate approach, Dynamic Feature Sharing (DFS) [128] identified relevant feature data for sharing through Random-K priorities based on Manhattan distance. DFS incorporated early, intermediate, and late fusion strategies within the CP system to optimize dynamic node engagement.
@@
While distance-based methods emphasize the transmission of feature cells or data points close to the ego vehicle, they may only partially meet the ego vehicles’ perception information requirements in real-world scenarios. In contrast, learning-based approaches leverage the entirety of perception data to intelligently determine information selection for sharing. In BM2CP [129], a modality-guided collaboration was designed to selectively share the most critical multi-modal features across agents, thereby promoting communication efficiency. This was achieved by generating a preference threshold mask to filter BEV features. Deep Reinforcement Learning (DRL) allows agents to learn policies through interactions with their environment, maximizing cumulative rewards [156, 157, 158]. The work [130] introduced a DRL-based Cooperative Perception (DRLCP) scheme, where each CAV intelligently decides what information to transmit based on the context captured by its onboard sensors. DRLCP’s primary objective is to determine whether to transmit or discard data, but it was evaluated on only two road networks without considering the impact of obstructions on vehicular communications.
@@
Another category of learning-based approaches, the selective-region mechanism, focuses on selecting CPM information within appropriate communication regions. For instance, [131] employed a bounding box matching module and strategic critical point selection to choose features enclosed within bounding box proposals for transmission. In contrast, UMC [132] introduced the Entropy-CS, a trainable region-wise communication selection module, leveraging entropy to distinguish informative regions and select suitable regions for transmission based on resolution levels. GevBEV [133] used evidential BEV maps to identify areas where the ego vehicle needs additional information, applying evidential deep learning with Dirichlet evidence [159] to quantify point-based classification uncertainty, thereby highlighting the most relevant regions for communication.
@@
Foreground regions contain more informative content than background regions. Based on this principle, the IFTR [134] employed a message selection and feature map reconstruction module to share the most relevant visual features from the foreground regions. In CodeFilling [135], the information-filling-driven message selection was designed to determine the ego vehicle’s perceptual needs by solving a local optimization problem based on the information score map gathered from other agents. An optimized selection matrix is produced to dictate which spatial regions should be prioritized for transmission. Spatial selection transmits only the informative regions identified by spatial confidence maps, but as bandwidth constraints become more stringent, this approach may lead to the potential loss of objects [136]. In contrast, V2X-PC [136] introduced a Point Cluster Packing module to control bandwidth usage by sampling important points. The point cluster representation preserves object features during message packing, enables efficient message aggregation regardless of collaboration range, and allows for explicit communication of object structure information.
@@
Attention mechanism enables a single agent to intelligently identify the most relevant perception information from multiple agents. For example, the pillar attention encoder (PAE) [137] extracted the attention value of each feature as an importance indicator and selected the most important features for transmission based on communication bandwidth. In ActFormer [138], an active selection network was proposed to assign each query with an interest score, which indicates the relevance and significance of each feature for the ego’s perception. This method actively selects the most relevant sensory data from multiple robots based on spatial knowledge without relying on the sensory measurements themselves, which improves the efficiency of multi-agent object detection.
@@ IV-B3 Agent and Information Selection
Methods that consider both cooperative agent and information selection jointly can offer a more balanced approach to optimizing precision and bandwidth in CP. One such approach, Where2comm [139], introduces a spatial-confidence-aware CP framework that employs spatial confidence maps generated by each agent to identify perceptually critical regions in the feature map. These critical features are compactly packaged into messages and shared via a sparsely connected communication graph. A communication graph construction strategy is employed to selectively determine communication partners, thereby minimizing unnecessary bandwidth usage. Furthermore, only non-zero features and their corresponding indices are transmitted. Experimental results indicate that Where2comm effectively facilitates mutually beneficial cooperation by sharing spatially sparse yet crucial features through V2X communication, optimizing both perception accuracy and communication efficiency.
@@ IV-B4 Perception Information Compression
Compression techniques can effectively prioritize and retain the most critical features of perception data, ensuring that even with reduced data size, the essential information required for perception tasks is preserved. For example, V2VNet [140] utilized a variational image compression algorithm [160] to compress feature maps. In [141], the feature maps from each vehicle are compressed through quantization to reduce communication costs. This on-demand sharing mechanism enabled efficient cooperation among connected vehicles by selectively sharing compressed feature maps but ignored the integration of spatial-temporal information. CPSC [142] performed feature-level compression by focusing on critical regions of perceptual information in the spatial-temporal domain while adapting the data transmission strategy based on network conditions.
@@
Feature encoding using Convolutional Neural Networks (CNNs) is a widely adopted method for feature compression. FS-COD [143] employed a filter at the last layer of the CNN-based feature extractor to control the size of the shared data between cooperative vehicles. DiscoNet [144] used a $1\times 1$ convolutional autoencoder [161] to compress feature maps along the channel dimension, while V2X-ViT [145] applied multiple $1\times 1$ convolutional filters for the same purpose. COOPERNAUT [146] introduced a Point Encoder using the Point Transformer [162] to compress point clouds into compact representations, filtering partially processed data to balance perception accuracy and communication load. Similarly, CenterCoop [147] encoded informative cues from the entire BEV context into compact center representations, significantly reducing communication costs by transmitting these compact center features as queries instead of the entire feature map.
@@
While CNN-generated feature maps offer potential data volume reduction, their dimensions often exceed transmission feasibility within current V2X technology scopes [148]. In [148], the SENet channel attention module [163] was used to weigh channel importance, determining optimal channels for transmission based on attention weights and the uniqueness of semantic information. EMIFF [149] introduced a Feature Compression module with channel and spatial compression blocks for transmission efficiency. Although these approaches effectively reduced the necessary transmission bandwidth, they did not account for the heterogeneity of transmitted information resulting from configuration disparities among agents [164]. To address this issue, What2comm [150] introduced a communication mechanism based on feature decoupling that strategically minimizes bandwidth usage by eliminating background noise to isolate sparse yet essential exclusive features and by compressing common features along the channel dimension.
@@
The aforementioned compression methods have uniformly applied the same compression ratio to all CAV data, but this fairness approach fails to account for the varying contributions of CAVs to perception during data fusion [165]. SmartCooper [151] optimized vehicle connectivity while considering communication constraints, using a learnable encoder to dynamically adjust the compression ratio based on channel state information and employing a judger mechanism to filter out detrimental image data.

### v5

@@ III Development of Cooperative Perception: From Early Time Exploration till Recent Research
Early research in CP focused on rule-based and statistical signal processing methods, often relying on radar- or LiDAR-generated data or artificial sources. These approaches employed classical tracking and motion models (e.g., particle filters, Kalman filters, GM-PHD filters) to align and fuse object-level data from multiple vehicles [16, 17, 18, 19]. While these methods demonstrated engineering viability and improved perception coverage, they were limited by rigid motion assumptions, sensor noise sensitivity, and lack of scalability to real-world traffic dynamics. Additional efforts incorporated biologically inspired algorithms, such as mirror-neuron models for intention awareness [20], though these approaches remained dependent on accurate pose estimation and often struggled in dynamic, multi-agent environments.
As CP evolved, limitations in extracting semantically rich and generalizable representations from shared data became apparent. This led to growing interest in deep learning, which has transformed single-vehicle perception and inspired learning-based CP frameworks to address traditional shortcomings. For example, cooperative object detection and classification tasks now benefit from deep architectures like YOLO and DenseNet to process shared data more effectively [21, 22, 23]. Recent developments in this direction are discussed further in Sections IV and V.
Historically, CP systems have relied on late collaboration due to feature extraction complexity and communication constraints. With advances in deep learning, intermediate collaboration is gaining traction by enabling efficient feature sharing among agents [24]. However, this approach introduces challenges such as information redundancy and the need for robust data alignment and fusion algorithms. Optimizing the relevance of shared features and selecting appropriate collaborators are critical for effective multi-agent cooperation. Consequently, developing unified CP frameworks that balance communication efficiency, perception accuracy, and scalability is key to advancing Cooperative Driving Automation (CDA).
@@ IV-A Framework Overview
TABLE II: Summary of the Methods of Preparation for Transmission.
@@
CAS Who2com [28] 2020 - Camera AirSim-CP Intermediate Segmentation $\surd$
When2com [29] 2020 - Camera AirSim-MAP Intermediate Segmentation $\surd$
@@
PACP [36] 2024 V2V Camera OPV2V Intermediate Detection $\times$
PIS DyDisse [37] 2019 V2X Camera/LiDAR - Late Tracking $\times$
VehIdenti [38] 2023 V2X - CARTERY - - $\times$
RedunMitig [39] 2019 V2X - SUMO Late Detection $\times$
Look-ahead [40] 2019 V2X - SUMO Late Detection $\times$
AICP [41] 2022 V2V Camera - Late Detection,Tracking $\surd$
@@
DRLCP [44] 2020 V2X Camera CIVS Late - $\times$
@@ IV-B Preparation for Transmission
In CP systems, continuous information exchange among multiple agents is crucial for achieving comprehensive traffic awareness. However, real-world application scenarios often face limitations in communication bandwidth. Excessive network load increases the risk of important data packets being delayed or lost, which could compromise the performance and safety of CP systems. Therefore, it is essential to balance the trade-off between perception performance and communication efficiency. To address this challenge, researchers have proposed various strategies aimed at reducing the volume of shared data prior to transmission and alleviating the burden on vehicular networks. As summarized in Table II, these strategies can be grouped into four main approaches:
@@ IV-B1 Cooperative Agent Selection
The attention mechanism, recognized for its efficacy in improving model efficiency by focusing on relevant data, offers promise in selecting suitable agents for the ego CAV. In this context, a multi-stage handshake communication mechanism was proposed in Who2com [28], employing the general attention mechanism to select the appropriate collaborators. This mechanism computes matching scores among agents, selectively engaging the most pertinent agents to optimize bandwidth usage. Building upon this, a scaled general attention mechanism was introduced in When2com [29] to determine optimal communication timing. Through this mechanism, communication was initiated by the ego vehicle only when its own information was deemed insufficient, thereby allowing network resources to be conserved. Although this multi-stage communication process proves effective, its implementation and optimization in dynamically varying environments pose significant challenges. To address this, a dynamic communication mechanism was proposed in [30], allowing vehicles to determine, based on learned attention weights, which infrastructure nodes to engage with, thereby balancing detection accuracy and communication resource utilization.
@@
The selection of cooperative agents can also be likened to the Multi-agent Path Finding (MAPF) problem, in which paths are planned for multiple agents [68]. Inspired by MAPF, a selective communication algorithm was proposed in [33] to determine a subset of cooperative vehicles based on estimated information gains from connected vehicles. In this approach, information gains were estimated by the ego vehicle through comparison between its own detected object positions and the received two-dimensional center positions from other connected vehicles. Vehicles associated with the highest estimated information gains were then selected to define the communication scope. In contrast to random selection strategies, this algorithm demonstrated higher success rates while incurring minimal additional communication costs.
Addressing the need for prioritization in CP, the Priority-Aware Collaborative Perception (PACP) framework leverages a Bird’s Eye View (BEV) based matching strategy to determine the relative importance of nearby CAVs based on their spatial correlation with the ego vehicle [36].
All the aforementioned methods primarily focused on spatial correlations, overlooking the significance of temporal dependencies in collaborator selection. To address this limitation, IoSI-CP [34] was proposed to exploit semantic information across both temporal and spatial dimensions for optimal collaborator selection. This approach leveraged a Graph Neural Network (GNN) to capture inherent correlations between the ego vehicle and its neighbouring agents. Agents with weights of zero or lower were considered ineligible for collaboration, which eliminates unsuitable collaborators responsible for noise. Nevertheless, this approach relied on one-dimensional weights that may not fully capture the intricacies addressed by high-dimensional representation methods. To enhance collaborator relevance, InterCoop [35] was developed to compute an interaction score for each vehicle by analyzing geometric properties of the road network in conjunction with the temporal dynamics of trajectories, thereby prioritizing vehicles most relevant to the prevailing driving conditions.
@@ IV-B2 Perception Information Selection
The rule-based approach to CPM generation/selection entails designing rules that dictate which aspects of detected object information should be included in a CPM. For example, in [37], a rule was proposed in which information about an object’s mode is transmitted only if the object is turning or accelerating. Similarly, in [38], visual information from detected vehicles was used to determine their communication identifiers, which were then employed to integrate CPMs with local perception data. All detected vehicles that have exchanged CAM messages will not be considered in future CPMs. Standardization efforts, led by the European Telecommunications Standards Institute (ETSI), have also introduced CPM generation rules, specifying what information CPMs should include and when to generate new messages. In [39] and [40], improved algorithms were proposed to refine these rules by filtering out redundant information about detected objects that had been recently transmitted by nearby vehicles. Although rule-based methods are relatively straightforward to implement, they primarily rely on changes in the position and speed of detected objects.
@@
Distance-based methods prioritize data based on distance metrics, considering the spatial location relative to the ego vehicle’s onboard sensors to measure data point significance. For instance, the Augmented Informative Collaborative Perception (AICP) system [41] employed a fitness sorting algorithm based on Mahalanobis distance to select crucial information at the application level, effectively filtering out less relevant packets with minimal delay. In a separate approach, Dynamic Feature Sharing (DFS) [42] was developed to identify relevant feature data for sharing using Random-K priorities based on Manhattan distance. Within this framework, early, intermediate, and late fusion strategies were incorporated to optimize dynamic node engagement in the cooperative perception system.
@@
While distance-based methods emphasize the transmission of feature cells or data points close to the ego vehicle, they may only partially meet the ego vehicles’ perception information requirements in real-world scenarios. In contrast, learning-based approaches leverage the entirety of perception data to intelligently determine information selection for sharing. In BM2CP [43], a modality-guided collaboration was designed to selectively share the most critical multi-modal features across agents, thereby promoting communication efficiency. This was achieved by generating a preference threshold mask to filter BEV features. Deep Reinforcement Learning (DRL) allows agents to learn policies through interactions with their environment, maximizing cumulative rewards [69]. The work [44] introduced a DRL-based Cooperative Perception (DRLCP) scheme, where each CAV intelligently decides what information to transmit based on the context captured by its onboard sensors. DRLCP’s primary objective is to determine whether to transmit or discard data, but it was evaluated on only two road networks without considering the impact of obstructions on vehicular communications.
@@
Another category of learning-based approaches, the selective-region mechanism, focuses on selecting CPM information within appropriate communication regions. For instance, in [45], a bounding box matching module was employed, and strategic critical point selection was adopted to choose features enclosed within bounding box proposals for transmission. In contrast, UMC [46] was proposed with a trainable region-wise communication selection module, Entropy-CS, in which entropy was leveraged to distinguish informative regions and select suitable areas for transmission based on resolution levels. GevBEV [47] was developed to use evidential BEV maps for identifying areas where the ego vehicle required additional information, and evidential deep learning with Dirichlet evidence [70] was applied to quantify point-based classification uncertainty, thereby highlighting the most relevant regions for communication.
@@
Foreground regions contain more informative content than background regions. Based on this principle, a message selection and feature map reconstruction module was employed in IFTR [48] to transmit the most relevant visual features from foreground regions. In CodeFilling [49], the information-filling-driven message selection was designed to determine the ego vehicle’s perceptual needs by solving a local optimization problem based on the information score map gathered from other agents. An optimized selection matrix was generated to determine which spatial regions should be prioritized for transmission. In spatial selection approaches, only the informative regions identified by spatial confidence maps are transmitted; however, under strict bandwidth constraints, this may result in the loss of critical object information [50]. To address this issue, a Point Cluster Packing module was introduced in V2X-PC [50], in which important points were sampled to control bandwidth usage. This point cluster representation was shown to preserve object features during message packing, support efficient message aggregation regardless of collaboration range, and enable the explicit communication of object structural information.
@@
Attention mechanism enables a single agent to intelligently identify the most relevant perception information from multiple agents. For example, in the pillar attention encoder (PAE) [51], attention values were extracted for each feature to serve as importance indicators, and the most significant features were selected for transmission based on available communication bandwidth. In ActFormer [52], an active selection network was proposed to assign each query with an interest score, which indicates the relevance and significance of each feature for the ego’s perception. This method actively selects the most relevant sensory data from multiple robots based on spatial knowledge without relying on the sensory measurements themselves, which improves the efficiency of multi-agent object detection.
@@ IV-B3 Agent and Information Selection
Methods that consider both cooperative agents and information selection jointly can offer a more balanced approach to optimizing precision and bandwidth in CP. One such method, Where2comm [53], was introduced as a spatial-confidence-aware cooperative perception framework in which spatial confidence maps are generated by each agent to identify perceptually critical regions in the feature map. These critical features were compactly packaged into messages and shared via a sparsely connected communication graph. A communication graph construction strategy was employed to selectively determine communication partners, thereby minimizing unnecessary bandwidth usage. Furthermore, only non-zero features and their corresponding indices were transmitted. Experimental results have demonstrated that Where2comm enables mutually beneficial cooperation by transmitting spatially sparse yet crucial features through vehicle-to-everything communication, thereby optimizing both perception accuracy and communication efficiency.
@@ IV-B4 Perception Information Compression
Compression techniques have been shown to effectively prioritize and retain the most critical features of perception data, ensuring that essential information required for perception tasks is preserved even when data size is reduced. For example, in V2VNet [54], a variational image compression algorithm [71] was employed to compress feature maps. In [55], feature maps from each vehicle were compressed through quantization to reduce communication costs. This on-demand sharing mechanism enabled efficient cooperation among connected vehicles by selectively transmitting compressed feature maps; however, the integration of spatial-temporal information was not considered. In CPSC [56], feature-level compression was performed by focusing on critical regions of perceptual information in the spatial-temporal domain, and the data transmission strategy was adapted based on network conditions.
@@
Feature encoding using Convolutional Neural Networks (CNNs) has been widely adopted for feature compression in cooperative perception. In FS-COD [57], a filter was applied at the final layer of the CNN-based feature extractor to control the size of the data shared between cooperative vehicles. In DiscoNet [58], a $1\times 1$ convolutional autoencoder [72] was utilized to compress feature maps along the channel dimension, while in V2X-ViT [59], multiple $1\times 1$ convolutional filters were applied for the same purpose. In COOPERNAUT [60], a Point Encoder based on the Point Transformer [73] was introduced to compress point clouds into compact representations, enabling the filtering of partially processed data to balance perception accuracy and communication load. Similarly, in CenterCoop [61], informative cues from the full bird’s-eye view context were encoded into compact center representations, and communication costs were significantly reduced by transmitting these compact features as queries instead of the full feature map.
@@
While CNN-generated feature maps offer potential data volume reduction, their dimensions often exceed transmission feasibility within current V2X technology scopes [62]. In [62], the SENet channel attention module [74] was used to weigh channel importance, determining optimal channels for transmission based on attention weights and the uniqueness of semantic information. In EMIFF [63], a Feature Compression module was introduced, incorporating both channel and spatial compression blocks to enhance transmission efficiency. Although these approaches effectively reduced the required transmission bandwidth, the heterogeneity of transmitted information—caused by configuration disparities among agents—was not addressed [75]. To address this issue, a communication mechanism based on feature decoupling was introduced in What2comm [64], in which bandwidth usage was strategically minimized by eliminating background noise to isolate sparse yet essential exclusive features, and by compressing common features along the channel dimension.
@@
The aforementioned compression methods have uniformly applied the same compression ratio to all CAV data. However, this fairness approach fails to account for the varying contributions of CAVs to perception during data fusion [76]. To address this limitation, SmartCooper [65] was proposed to optimize vehicle connectivity under communication constraints. In this method, a learnable encoder was used to dynamically adjust the compression ratio based on channel state information, and a judger mechanism was employed to filter out detrimental image data.